# Natural Language Generation from Visual Events: Challenges and Future Directions

### v2

@@ 6 Open Problems and Research Directions
Many of the tasks we consider in this work have real-world applications. For instance, solutions to the Change Captioning task can be used for assisted surveillance and for tracking changes in digital media assets Jhamtani and Berg-Kirkpatrick (2018). In the Movie Auto AD task, models are required to generate descriptions that complement information in the original audio dialog and soundtrack, for improved accessibility to visually impaired users and for enhancing the visual experience of sighted users Han et al. (2023a).
However, many day-to-day human-centered scenarios involve personalizing to various contexts or situations. We argue that existing tasks in this domain, in their definitions and settings, do not fully reflect this aspect. Tailoring model-generated descriptions or narrations to the perspective of end-users requires task settings in which models and humans can interact iteratively. Such settings would enable the incorporation of human expectations and communicative contexts, which typically tend to be dynamic in real-world applications. To this end, we advocate for variations of existing tasks where models can learn to contextualize and reason through interactions with other agents (humans or other models) for generating stories, descriptions, or answers. In vision-to-language tasks involving singular images, various works have explored controlled settings in which models are expected to generate text adhering to a specific style Tan et al. (2022); Wang et al. (2023) or user-preferences Alaluf et al. (2025); Pi et al. (2025). Models developed for solving these variations of tasks can be directly deployed to meet the requirements of real-world use cases. However, in tasks that require generating text from multiple images or video frames, there has been limited work concerning the abilities of models to control for the style of the generated output Yang and Jin (2023). We advocate for an exploration of such settings in language generation tasks dealing with visual events, for enabling the development of models that can adapt and cater to various points of view or genres dynamically.
@@
As discussed in Section 2.4, VLMs that are trained on various general-purpose datasets are increasingly being used in these scenarios through prompting. Powerful open-source models such as Molmo Deitke et al. (2024), and models optimized for multi-image scenarios such as Mantis Jiang et al. (2024) are becoming increasingly available, suggesting that the trend of adopting them off-the-shelf for solving many vision-to-language tasks is widespread. General-purpose VLMs learn abundant information through multitask pre-training and have a modular design, making them suitable for many downstream tasks. Their modularity also enables seamless adaptation of VLMs to various novel domains (e.g., medical science) by updating only a small fraction of their parameters Li et al. (2023a).
Despite the promising generalization of VLMs to certain tasks and domains, they have also been shown to be sensitive to prompts Liu et al. (2024); Schlarmann and Hein (2023) and biased towards the textual modality Rahmanzadehgervi et al. (2024). To address these problems, recent work proposes various prompt engineering techniques to facilitate inference-time adaptation of prompts to make them more suitable for the specific task of interest Ma et al. (2023); Gu et al. (2023a). On the other hand, task-specific model architectures consist of components designed to effectively address specialized aspects of the tasks, e.g., computing a difference representation of the input image pair in Change Captioning. We advocate for modular modeling approaches that bring together efficient task-specific components and combine them with the powerful foundational VLMs. Furthermore, we argue that using graph-based architectures and memory-based modules akin to Zhang et al. (2024b); Feng et al. (2024) would result in improved tracking of entity positions and relationships, and enable models to assign saliency to memorable events in tasks like Visual Storytelling or Movie Auto AD.
@@
In Section 3, we discussed various approaches for evaluating model outputs in this domain. While human evaluation is impractical for conducting large-scale assessments, existing automatic evaluation metrics are limited in terms of fully capturing the abilities of models. Increasingly, various benchmarking datasets are being proposed to assess models along different axes important for grounding language in visual events Li et al. (2024b). However, many benchmarks often suffer from the problem of visual content irrelevance, which refers to models performing well on the benchmark datasets by primarily relying only on the language modality Chen et al. (2024a). Furthermore, data leakage and contamination problems (see section 3.1) also hinder fair and accurate testing of a model’s skills using benchmarks.
@@
While it is important to continue directing research efforts towards developing more extensive multi-image benchmarks such as ReMI Kazemi et al. (2024), we argue that the purpose of evaluation is to also provide insights that can be directly leveraged for improving model architectures and learning procedures. For vision-to-language tasks involving single images, such as visual question answering and image captioning, recent work has used various interpretability methods for obtaining insights about the relations between input-output data Cafagna et al. (2023); Parcalabescu and Frank (2023) and for explaining model behavior Wang, Rudner, and Wilson (2023); Neo et al. (2024); Peng et al. (2024). There has also been work that focuses on understanding the functioning of general-purpose VLMs by analyzing the internal mechanisms and interactions between their components Yu and Ananiadou (2024); Huo et al. (2024). These methods can complement traditional evaluation techniques for enabling intra- and inter-model comparisons to holistically understand behaviors and representations required for conditioning language in the visual data Lin et al. (2025). For scenarios dealing with language generation from visual events, interpretability methods are largely unexplored. We strongly advocate for efforts to develop interpretability techniques tailored for NLG tasks dealing with visual events, and also for adapting existing approaches from unimodal and vision-to-language settings involving static images.
7 Conclusion
@@
In this position paper, we focus on the problem of generating language from visual events and gather, under the same umbrella, various tasks that are typically considered separate by the research community. Despite having different characteristics in terms of their communicative intents or input-output data, we argued that all the tasks present a common set of challenges for developing models and assessing their generated outputs. To understand the progress made over the years, we extensively reviewed the different modeling approaches and evaluation protocols for each of the various exemplificative tasks we considered, and discussed them in a comprehensive and unified manner. We also proposed a method for quantitatively exploring the current landscape of this problem, using which we capture differences between the tasks in terms of their visual input and language output. As the problem of generating text conditioned on sequences of multiple temporally ordered images or video frames has various real-world applications, we underline the challenges and propose several concrete research directions informed by insights from linguistics, cognitive sciences, and natural language processing (NLP) aimed towards facilitating further advancements. We argue that leveraging these insights could also help the development of better VLMs, which are currently not immune from some of the highlighted limitations.

### v3

@@ 6. Open Problems and Research Directions
Many of the tasks we consider in this work have real-world applications. For instance, solutions to the Change Captioning task can be used for assisted surveillance and for tracking changes in digital media assets (Jhamtani and Berg-Kirkpatrick, 2018). In the Movie Auto AD task, models are required to generate descriptions that complement information in the original audio dialog and soundtrack, for improved accessibility to visually impaired users and for enhancing the visual experience of sighted users (Han et al., 2023a).
However, many day-to-day human-centered scenarios involve personalizing to various contexts or situations. We argue that existing tasks in this domain, in their definitions and settings, do not fully reflect this aspect. Tailoring model-generated descriptions or narrations to the perspective of end-users requires task settings in which models and humans can interact iteratively. Such settings would enable the incorporation of human expectations and communicative contexts, which typically tend to be dynamic in real-world applications. To this end, we advocate for variations of existing tasks where models can learn to contextualize and reason through interactions with other agents (humans or other models) for generating stories, descriptions, or answers. In vision-to-language tasks involving singular images, various works have explored controlled settings in which models are expected to generate text adhering to a specific style (Tan et al., 2022; Wang et al., 2023b) or user-preferences (Alaluf et al., 2025; Pi et al., 2025). Models developed for solving these variations of tasks can be directly deployed to meet the requirements of real-world use cases. However, in tasks that require generating text from multiple images or video frames, there has been limited work concerning the abilities of models to control for the style of the generated output (Yang and Jin, 2023). We advocate for an exploration of such settings in language generation tasks dealing with visual events, for enabling the development of models that can adapt and cater to various points of view or genres dynamically.
@@
As discussed in Section 2.4, VLMs that are trained on various general-purpose datasets are increasingly being used in these scenarios through prompting. Powerful open-source models such as Molmo (Deitke et al., 2024), and models optimized for multi-image scenarios such as Mantis (Jiang et al., 2024) are becoming increasingly available, suggesting that the trend of adopting them off-the-shelf for solving many vision-to-language tasks is widespread. General-purpose VLMs learn abundant information through multitask pre-training and have a modular design, making them suitable for many downstream tasks. Their modularity also enables seamless adaptation of VLMs to various novel domains (e.g., medical science) by updating only a small fraction of their parameters (Li et al., 2023b).
Despite the promising generalization of VLMs to certain tasks and domains, they have also been shown to be sensitive to prompts (Liu et al., 2024; Schlarmann and Hein, 2023) and biased towards the textual modality (Rahmanzadehgervi et al., 2024). To address these problems, recent work proposes various prompt engineering techniques to facilitate inference-time adaptation of prompts to make them more suitable for the specific task of interest (Ma et al., 2023; Gu et al., 2023b). On the other hand, task-specific model architectures consist of components designed to effectively address specialized aspects of the tasks, e.g., computing a difference representation of the input image pair in Change Captioning. We advocate for modular modeling approaches that bring together efficient task-specific components and combine them with the powerful foundational VLMs. Furthermore, we argue that using graph-based architectures and memory-based modules akin to Zhang et al. (2024b); Feng et al. (2024) would result in improved tracking of entity positions and relationships, and enable models to assign saliency to memorable events in tasks like Visual Storytelling or Movie Auto AD.
@@
In Section 4, we discussed various approaches for evaluating model outputs in this domain. While human evaluation is impractical for conducting large-scale assessments, existing automatic evaluation metrics are limited in terms of fully capturing the abilities of models. Increasingly, various benchmarking datasets are being proposed to assess models along different axes important for grounding language in visual events (Li et al., 2024a). However, many benchmarks often suffer from the problem of visual content irrelevance, which refers to models performing well on the benchmark datasets by primarily relying only on the language modality (Chen et al., 2024a). Furthermore, data leakage and contamination problems (see section 4.1) also hinder fair and accurate testing of a model’s skills using benchmarks.
@@
While it is important to continue directing research efforts towards developing more extensive multi-image benchmarks such as ReMI (Kazemi et al., 2024), we argue that the purpose of evaluation is to also provide insights that can be directly leveraged for improving model architectures and learning procedures. For vision-to-language tasks involving single images, such as visual question answering and image captioning, recent work has used various interpretability methods for obtaining insights about relations between the visual and textual modalities of data (Cafagna et al., 2023; Parcalabescu and Frank, 2023), and for explaining model behavior (Wang et al., 2023a; Neo et al., 2024; Peng et al., 2024). There has also been work that focuses on understanding the functioning of general-purpose VLMs by analyzing the internal mechanisms and interactions between their components (Yu and Ananiadou, 2024; Huo et al., 2024). These methods can complement traditional evaluation techniques for enabling intra- and inter-model comparisons to holistically understand behaviors and representations required for conditioning language in the visual data (Lin et al., 2025). For scenarios dealing with language generation from visual events, interpretability methods are largely unexplored. We strongly advocate for efforts to develop interpretability techniques tailored for NLG tasks dealing with visual events, and also for adapting existing approaches from unimodal and vision-to-language settings involving static images.
7. Conclusion
@@
In this paper, we focused on the problem of generating language from visual events and gather, under the same umbrella, various tasks that are typically considered separate by the research community. Despite having different characteristics in terms of their communicative intents or input-output data, we argued that all the tasks present a common set of challenges for developing models and assessing their generated outputs. To understand the progress made over the years, we extensively reviewed the different modeling approaches and evaluation protocols for each of the various exemplificative tasks we considered, and discussed them in a comprehensive and unified manner. We also presented a small scale quantitative analysis in terms of the properties of datasets used to train models for the different tasks considered, regarding their visual and textual modalities. This type of data analysis can help to better understand the requirements of each task. The problem of generating text conditioned on sequences of multiple temporally ordered images or video frames has various real-world applications, which are becoming more present in our everyday lives. To facilitate further advancements, we underlined the challenges posed by this problem and propose several concrete research directions informed by insights from linguistics, cognitive sciences, and natural language processing. We argue that leveraging these insights could also help the development of better VLMs, which are currently not immune from some of the highlighted limitations.